# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ Acknowledgements.
- OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mohammad Bavarian, Jeff Belgum, Irwan Bello, Jake Berdine, Gabriel Bernadett-Shapiro, Christopher Berner, Lenny Bogdonoff, Oleg Boiko, Madelaine Boyd, Anna-Luisa Brakman, Greg Brockman, Tim Brooks, Miles Brundage, Kevin Button, Trevor Cai, Rosie Campbell, Andrew Cann, Brittany Carey, Chelsea Carlson, Rory Carmichael, Brooke Chan, Che Chang, Fotis Chantzis, Derek Chen, Sully Chen, Ruby Chen, Jason Chen, Mark Chen, Ben Chess, Chester Cho, Casey Chu, Hyung Won Chung, Dave Cummings, Jeremiah Currier, Yunxing Dai, Cory Decareaux, Thomas Degry, Noah Deutsch, Damien Deville, Arka Dhar, David Dohan, Steve Dowling, Sheila Dunning, Adrien Ecoffet, Atty Eleti, Tyna Eloundou, David Farhi, Liam Fedus, Niko Felix, Simón Posada Fishman, Juston Forte, Isabella Fulford, Leo Gao, Elie Georges, Christian Gibson, Vik Goel, Tarun Gogineni, Gabriel Goh, Rapha Gontijo-Lopes, Jonathan Gordon, Morgan Grafstein, Scott Gray, Ryan Greene, Joshua Gross, Shixiang Shane Gu, Yufei Guo, Chris Hallacy, Jesse Han, Jeff Harris, Yuchen He, Mike Heaton, Johannes Heidecke, Chris Hesse, Alan Hickey, Wade Hickey, Peter Hoeschele, Brandon Houghton, Kenny Hsu, Shengli Hu, Xin Hu, Joost Huizinga, Shantanu Jain, Shawn Jain, Joanne Jang, Angela Jiang, Roger Jiang, Haozhun Jin, Denny Jin, Shino Jomoto, Billie Jonn, Heewoo Jun, Tomer Kaftan, Łukasz Kaiser, Ali Kamali, Ingmar Kanitscheider, Nitish Shirish Keskar, Tabarak Khan, Logan Kilpatrick, Jong Wook Kim, Christina Kim, Yongjik Kim, Jan Hendrik Kirchner, Jamie Kiros, Matt Knight, Daniel Kokotajlo, Łukasz Kondraciuk, Andrew Kondrich, Aris Konstantinidis, Kyle Kosic, Gretchen Krueger, Vishal Kuo, Michael Lampe, Ikai Lan, Teddy Lee, Jan Leike, Jade Leung, Daniel Levy, Chak Ming Li, Rachel Lim, Molly Lin, Stephanie Lin, Mateusz Litwin, Theresa Lopez, Ryan Lowe, Patricia Lue, Anna Makanju, Kim Malfacini, Sam Manning, Todor Markov, Yaniv Markovski, Bianca Martin, Katie Mayer, Andrew Mayne, Bob McGrew, Scott Mayer McKinney, Christine McLeavey, Paul McMillan, Jake McNeil, David Medina, Aalok Mehta, Jacob Menick, Luke Metz, Andrey Mishchenko, Pamela Mishkin, Vinnie Monaco, Evan Morikawa, Daniel Mossing, Tong Mu, Mira Murati, Oleg Murk, David Mély, Ashvin Nair, Reiichiro Nakano, Rajeev Nayak, Arvind Neelakantan, Richard Ngo, Hyeonwoo Noh, Long Ouyang, Cullen O’Keefe, Jakub Pachocki, Alex Paino, Joe Palermo, Ashley Pantuliano, Giambattista Parascandolo, Joel Parish, Emy Parparita, Alex Passos, Mikhail Pavlov, Andrew Peng, Adam Perelman, Filipe de Avila Belbute Peres, Michael Petrov, Henrique Ponde de Oliveira Pinto, Michael, Pokorny, Michelle Pokrass, Vitchyr H. Pong, Tolly Powell, Alethea Power, Boris Power, Elizabeth Proehl, Raul Puri, Alec Radford, Jack Rae, Aditya Ramesh, Cameron Raymond, Francis Real, Kendra Rimbach, Carl Ross, Bob Rotsted, Henri Roussez, Nick Ryder, Mario Saltarelli, Ted Sanders, Shibani Santurkar, Girish Sastry, Heather Schmidt, David Schnurr, John Schulman, Daniel Selsam, Kyla Sheppard, Toki Sherbakov, Jessica Shieh, Sarah Shoker, Pranav Shyam, Szymon Sidor, Eric Sigler, Maddie Simens, Jordan Sitkin, Katarina Slama, Ian Sohl, Benjamin Sokolowsky, Yang Song, Natalie Staudacher, Felipe Petroski Such, Natalie Summers, Ilya Sutskever, Jie Tang, Nikolas Tezak, Madeleine B. Thompson, Phil Tillet, Amin Tootoonchian, Elizabeth Tseng, Preston Tuggle, Nick Turley, Jerry Tworek, Juan Felipe Cerón Uribe, Andrea Vallone, Arun Vijayvergiya, Chelsea Voss, Carroll Wainwright, Justin Jay Wang, Alvin Wang, Ben Wang, Jonathan Ward, Jason Wei, C. J. Weinmann, Akila Welihinda, Peter Welinder, Jiayi Weng, Lilian Weng, Matt Wiethoff, Dave Willner, Clemens Winter, Samuel Wolrich, Hannah Wong, Lauren Workman, Sherwin Wu, Jeff Wu, Michael Wu, Kai Xiao, Tao Xu, Sarah Yoo, Kevin Yu, Qiming Yuan, Wojciech Zaremba, Rowan Zellers, Chong Zhang, Marvin Zhang, Shengjia Zhao, Tianhao Zheng, Juntang Zhuang, William Zhuk, and Barret Zoph. 2024. GPT-4 Technical Report. doi:10.48550/arXiv.2303.08774 arXiv:2303.08774 [cs].
@@
- Pires de Lima et al. (2023) Rafael Pires de Lima, Behzad Vahedi, Nick Hughes, Andrew P Barrett, Walter Meier, and Morteza Karimzadeh. 2023. Enhancing sea ice segmentation in Sentinel-1 images with atrous convolutions. International Journal of Remote Sensing 44, 17 (2023), 5344–5374.
@@
- Poursanidis et al. (2019) Dimitris Poursanidis, Dimosthenis Traganos, Peter Reinartz, and Nektarios Chrysoulakis. 2019. On the use of Sentinel-2 for coastal habitat mapping and satellite-derived bathymetry estimation using downscaled coastal aerosol band. International Journal of Applied Earth Observation and Geoinformation 80 (Aug. 2019), 58–70. doi:10.1016/j.jag.2019.03.012
@@
- Press et al. (2022) Ofir Press, Noah A. Smith, and Mike Lewis. 2022. Train Short, Test Long: Attention with Linear Biases Enables Input Length Extrapolation. doi:10.48550/arXiv.2108.12409 arXiv:2108.12409 [cs].
@@
- Reed et al. (2023) Colorado J. Reed, Ritwik Gupta, Shufan Li, Sarah Brockman, Christopher Funk, Brian Clipp, Kurt Keutzer, Salvatore Candido, Matt Uyttendaele, and Trevor Darrell. 2023. Scale-MAE: A Scale-Aware Masked Autoencoder for Multiscale Geospatial Representation Learning. doi:10.48550/arXiv.2212.14532 arXiv:2212.14532 [cs].
@@
- Rolf et al. (2024) Esther Rolf, Konstantin Klemmer, Caleb Robinson, and Hannah Kerner. 2024. Mission Critical – Satellite Data is a Distinct Modality in Machine Learning. doi:10.48550/arXiv.2402.01444 arXiv:2402.01444 [cs].
@@
- Rolf et al. (2021) Esther Rolf, Jonathan Proctor, Tamma Carleton, Ian Bolliger, Vaishaal Shankar, Miyabi Ishihara, Benjamin Recht, and Solomon Hsiang. 2021. A generalizable and accessible approach to machine learning with global satellite imagery. Nature Communications 12, 1 (July 2021), 4392. doi:10.1038/s41467-021-24638-z
@@
- Rosenqvist et al. (2007) Ake Rosenqvist, Masanobu Shimada, Norimasa Ito, and Manabu Watanabe. 2007. ALOS PALSAR: A Pathfinder Mission for Global-Scale Monitoring of the Environment. IEEE Transactions on Geoscience and Remote Sensing 45, 11 (Nov. 2007), 3307–3316. doi:10.1109/TGRS.2007.901027 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Russakovsky et al. (2015) Olga Russakovsky, Jia Deng, Hao Su, Jonathan Krause, Sanjeev Satheesh, Sean Ma, Zhiheng Huang, Andrej Karpathy, Aditya Khosla, Michael Bernstein, Alexander C. Berg, and Li Fei-Fei. 2015. ImageNet Large Scale Visual Recognition Challenge. International Journal of Computer Vision 115, 3 (Dec. 2015), 211–252. doi:10.1007/s11263-015-0816-y
@@
- Rußwurm et al. (2023) Marc Rußwurm, Konstantin Klemmer, Esther Rolf, Robin Zbinden, and Devis Tuia. 2023. Geographic location encoding with spherical harmonics and sinusoidal representation networks. arXiv preprint arXiv:2310.06743 (2023).
@@
- Singh et al. (2022) Amanpreet Singh, Ronghang Hu, Vedanuj Goswami, Guillaume Couairon, Wojciech Galuba, Marcus Rohrbach, and Douwe Kiela. 2022. FLAVA: A Foundational Language And Vision Alignment Model. In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, New Orleans, LA, USA, 15617–15629. doi:10.1109/CVPR52688.2022.01519
@@
- Sitzmann et al. (2020) Vincent Sitzmann, Julien Martel, Alexander Bergman, David Lindell, and Gordon Wetzstein. 2020. Implicit neural representations with periodic activation functions. Advances in neural information processing systems 33 (2020), 7462–7473.
@@
- Solari et al. (2017) Lorenzo Solari, Andrea Ciampalini, Federico Raspini, Silvia Bianchini, Ivana Zinno, Manuela Bonano, Michele Manunta, Sandro Moretti, and Nicola Casagli. 2017. Combined Use of C- and X-Band SAR Data for Subsidence Monitoring in an Urban Area. Geosciences 7, 2 (June 2017), 21. doi:10.3390/geosciences7020021 Number: 2 Publisher: Multidisciplinary Digital Publishing Institute.
@@
- Spoto et al. (2012) Francois Spoto, Omar Sy, Paolo Laberinti, Philippe Martimort, Valerie Fernandez, Olivier Colin, Bianca Hoersch, and Aime Meygret. 2012. Overview Of Sentinel-2. In 2012 IEEE International Geoscience and Remote Sensing Symposium. 1707–1710. doi:10.1109/IGARSS.2012.6351195 ISSN: 2153-7003.
@@
- Sumbul et al. (2019) Gencer Sumbul, Marcela Charfuelan, Begüm Demir, and Volker Markl. 2019. Bigearthnet: A Large-Scale Benchmark Archive for Remote Sensing Image Understanding. In IGARSS 2019 - 2019 IEEE International Geoscience and Remote Sensing Symposium. 5901–5904. doi:10.1109/IGARSS.2019.8900532 ISSN: 2153-7003.
@@
- Sun et al. (2023) Xian Sun, Peijin Wang, Wanxuan Lu, Zicong Zhu, Xiaonan Lu, Qibin He, Junxi Li, Xuee Rong, Zhujun Yang, Hao Chang, Qinglin He, Guang Yang, Ruiping Wang, Jiwen Lu, and Kun Fu. 2023. RingMo: A Remote Sensing Foundation Model With Masked Image Modeling. IEEE Transactions on Geoscience and Remote Sensing 61 (2023), 1–22. doi:10.1109/TGRS.2022.3194732 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Tachi et al. (1989) K. Tachi, K. Arai, and Y. Sato. 1989. Advanced microwave scanning radiometer (AMSR): requirements and preliminary design study. IEEE Transactions on Geoscience and Remote Sensing 27, 2 (March 1989), 177–183. doi:10.1109/36.20296 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Thompson* (2015) Alan A. Thompson*. 2015. Overview of the RADARSAT Constellation Mission. Canadian Journal of Remote Sensing 41, 5 (Sept. 2015), 401–407. doi:10.1080/07038992.2015.1104633 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.1080/07038992.2015.1104633.
@@
- Torres et al. (2012) Ramon Torres, Paul Snoeij, Dirk Geudtner, David Bibby, Malcolm Davidson, Evert Attema, Pierre Potin, BjÖrn Rommen, Nicolas Floury, Mike Brown, Ignacio Navas Traver, Patrick Deghaye, Berthyl Duesmann, Betlem Rosich, Nuno Miranda, Claudio Bruno, Michelangelo L’Abbate, Renato Croci, Andrea Pietropaolo, Markus Huchler, and Friedhelm Rostan. 2012. GMES Sentinel-1 mission. Remote Sensing of Environment 120 (May 2012), 9–24. doi:10.1016/j.rse.2011.05.028
@@
- Tseng et al. (2024) Gabriel Tseng, Ruben Cartuyvels, Ivan Zvonkov, Mirali Purohit, David Rolnick, and Hannah Kerner. 2024. Lightweight, Pre-trained Transformers for Remote Sensing Timeseries. doi:10.48550/arXiv.2304.14065 arXiv:2304.14065 [cs].
@@
- Tsokas et al. (2022) Arsenios Tsokas, Maciej Rysz, Panos M. Pardalos, and Kathleen Dipple. 2022. SAR data applications in earth observation: An overview. Expert Systems with Applications 205 (Nov. 2022), 117342. doi:10.1016/j.eswa.2022.117342
- Vaswani et al. (2023) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2023. Attention Is All You Need. doi:10.48550/arXiv.1706.03762 arXiv:1706.03762 [cs].
@@
- Wang et al. (2023b) Di Wang, Qiming Zhang, Yufei Xu, Jing Zhang, Bo Du, Dacheng Tao, and Liangpei Zhang. 2023b. Advancing Plain Vision Transformer Toward Remote Sensing Foundation Model. IEEE Transactions on Geoscience and Remote Sensing 61 (2023), 1–15. doi:10.1109/TGRS.2022.3222818 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Wang et al. (2016) Qunming Wang, Wenzhong Shi, Zhongbin Li, and Peter M. Atkinson. 2016. Fusion of Sentinel-2 images. Remote Sensing of Environment 187 (Dec. 2016), 241–252. doi:10.1016/j.rse.2016.10.030

### v2

@@ Acknowledgements.
- OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mohammad Bavarian, Jeff Belgum, Irwan Bello, Jake Berdine, Gabriel Bernadett-Shapiro, Christopher Berner, Lenny Bogdonoff, Oleg Boiko, Madelaine Boyd, Anna-Luisa Brakman, Greg Brockman, Tim Brooks, Miles Brundage, Kevin Button, Trevor Cai, Rosie Campbell, Andrew Cann, Brittany Carey, Chelsea Carlson, Rory Carmichael, Brooke Chan, Che Chang, Fotis Chantzis, Derek Chen, Sully Chen, Ruby Chen, Jason Chen, Mark Chen, Ben Chess, Chester Cho, Casey Chu, Hyung Won Chung, Dave Cummings, Jeremiah Currier, Yunxing Dai, Cory Decareaux, Thomas Degry, Noah Deutsch, Damien Deville, Arka Dhar, David Dohan, Steve Dowling, Sheila Dunning, Adrien Ecoffet, Atty Eleti, Tyna Eloundou, David Farhi, Liam Fedus, Niko Felix, Simón Posada Fishman, Juston Forte, Isabella Fulford, Leo Gao, Elie Georges, Christian Gibson, Vik Goel, Tarun Gogineni, Gabriel Goh, Rapha Gontijo-Lopes, Jonathan Gordon, Morgan Grafstein, Scott Gray, Ryan Greene, Joshua Gross, Shixiang Shane Gu, Yufei Guo, Chris Hallacy, Jesse Han, Jeff Harris, Yuchen He, Mike Heaton, Johannes Heidecke, Chris Hesse, Alan Hickey, Wade Hickey, Peter Hoeschele, Brandon Houghton, Kenny Hsu, Shengli Hu, Xin Hu, Joost Huizinga, Shantanu Jain, Shawn Jain, Joanne Jang, Angela Jiang, Roger Jiang, Haozhun Jin, Denny Jin, Shino Jomoto, Billie Jonn, Heewoo Jun, Tomer Kaftan, Łukasz Kaiser, Ali Kamali, Ingmar Kanitscheider, Nitish Shirish Keskar, Tabarak Khan, Logan Kilpatrick, Jong Wook Kim, Christina Kim, Yongjik Kim, Jan Hendrik Kirchner, Jamie Kiros, Matt Knight, Daniel Kokotajlo, Łukasz Kondraciuk, Andrew Kondrich, Aris Konstantinidis, Kyle Kosic, Gretchen Krueger, Vishal Kuo, Michael Lampe, Ikai Lan, Teddy Lee, Jan Leike, Jade Leung, Daniel Levy, Chak Ming Li, Rachel Lim, Molly Lin, Stephanie Lin, Mateusz Litwin, Theresa Lopez, Ryan Lowe, Patricia Lue, Anna Makanju, Kim Malfacini, Sam Manning, Todor Markov, Yaniv Markovski, Bianca Martin, Katie Mayer, Andrew Mayne, Bob McGrew, Scott Mayer McKinney, Christine McLeavey, Paul McMillan, Jake McNeil, David Medina, Aalok Mehta, Jacob Menick, Luke Metz, Andrey Mishchenko, Pamela Mishkin, Vinnie Monaco, Evan Morikawa, Daniel Mossing, Tong Mu, Mira Murati, Oleg Murk, David Mély, Ashvin Nair, Reiichiro Nakano, Rajeev Nayak, Arvind Neelakantan, Richard Ngo, Hyeonwoo Noh, Long Ouyang, Cullen O’Keefe, Jakub Pachocki, Alex Paino, Joe Palermo, Ashley Pantuliano, Giambattista Parascandolo, Joel Parish, Emy Parparita, Alex Passos, Mikhail Pavlov, Andrew Peng, Adam Perelman, Filipe de Avila Belbute Peres, Michael Petrov, Henrique Ponde de Oliveira Pinto, Michael, Pokorny, Michelle Pokrass, Vitchyr H. Pong, Tolly Powell, Alethea Power, Boris Power, Elizabeth Proehl, Raul Puri, Alec Radford, Jack Rae, Aditya Ramesh, Cameron Raymond, Francis Real, Kendra Rimbach, Carl Ross, Bob Rotsted, Henri Roussez, Nick Ryder, Mario Saltarelli, Ted Sanders, Shibani Santurkar, Girish Sastry, Heather Schmidt, David Schnurr, John Schulman, Daniel Selsam, Kyla Sheppard, Toki Sherbakov, Jessica Shieh, Sarah Shoker, Pranav Shyam, Szymon Sidor, Eric Sigler, Maddie Simens, Jordan Sitkin, Katarina Slama, Ian Sohl, Benjamin Sokolowsky, Yang Song, Natalie Staudacher, Felipe Petroski Such, Natalie Summers, Ilya Sutskever, Jie Tang, Nikolas Tezak, Madeleine B. Thompson, Phil Tillet, Amin Tootoonchian, Elizabeth Tseng, Preston Tuggle, Nick Turley, Jerry Tworek, Juan Felipe Cerón Uribe, Andrea Vallone, Arun Vijayvergiya, Chelsea Voss, Carroll Wainwright, Justin Jay Wang, Alvin Wang, Ben Wang, Jonathan Ward, Jason Wei, C. J. Weinmann, Akila Welihinda, Peter Welinder, Jiayi Weng, Lilian Weng, Matt Wiethoff, Dave Willner, Clemens Winter, Samuel Wolrich, Hannah Wong, Lauren Workman, Sherwin Wu, Jeff Wu, Michael Wu, Kai Xiao, Tao Xu, Sarah Yoo, Kevin Yu, Qiming Yuan, Wojciech Zaremba, Rowan Zellers, Chong Zhang, Marvin Zhang, Shengjia Zhao, Tianhao Zheng, Juntang Zhuang, William Zhuk, and Barret Zoph. 2024. GPT-4 Technical Report. https://doi.org/10.48550/arXiv.2303.08774 arXiv:2303.08774 [cs].
- Oquab et al. (2024) Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, Mahmoud Assran, Nicolas Ballas, Wojciech Galuba, Russell Howes, Po-Yao Huang, Shang-Wen Li, Ishan Misra, Michael Rabbat, Vasu Sharma, Gabriel Synnaeve, Hu Xu, Hervé Jegou, Julien Mairal, Patrick Labatut, Armand Joulin, and Piotr Bojanowski. 2024. DINOv2: Learning Robust Visual Features without Supervision. https://doi.org/10.48550/arXiv.2304.07193 arXiv:2304.07193 [cs].
- Phuong and Lampert (2019) Mary Phuong and Christoph Lampert. 2019. Towards Understanding Knowledge Distillation. In Proceedings of the 36th International Conference on Machine Learning. PMLR, 5142–5151. https://proceedings.mlr.press/v97/phuong19a.html ISSN: 2640-3498.
@@
- Pires de Lima et al. (2023) Rafael Pires de Lima, Behzad Vahedi, Nick Hughes, Andrew P. Barrett, Walter Meier, and Morteza Karimzadeh. 2023. Enhancing sea ice segmentation in Sentinel-1 images with atrous convolutions. International Journal of Remote Sensing 44, 17 (Sept. 2023), 5344–5374. https://doi.org/10.1080/01431161.2023.2248560 Publisher: Taylor & Francis _eprint: https://doi.org/10.1080/01431161.2023.2248560.
@@
- Poursanidis et al. (2019) Dimitris Poursanidis, Dimosthenis Traganos, Peter Reinartz, and Nektarios Chrysoulakis. 2019. On the use of Sentinel-2 for coastal habitat mapping and satellite-derived bathymetry estimation using downscaled coastal aerosol band. International Journal of Applied Earth Observation and Geoinformation 80 (Aug. 2019), 58–70. https://doi.org/10.1016/j.jag.2019.03.012
@@
- Press et al. (2022) Ofir Press, Noah A. Smith, and Mike Lewis. 2022. Train Short, Test Long: Attention with Linear Biases Enables Input Length Extrapolation. https://doi.org/10.48550/arXiv.2108.12409 arXiv:2108.12409 [cs].
@@
- Radford et al. ([n. d.]) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. [n. d.]. Language Models are Unsupervised Multitask Learners. ([n. d.]).
@@
- Reed et al. (2023) Colorado J. Reed, Ritwik Gupta, Shufan Li, Sarah Brockman, Christopher Funk, Brian Clipp, Kurt Keutzer, Salvatore Candido, Matt Uyttendaele, and Trevor Darrell. 2023. Scale-MAE: A Scale-Aware Masked Autoencoder for Multiscale Geospatial Representation Learning. https://doi.org/10.48550/arXiv.2212.14532 arXiv:2212.14532 [cs].
@@
- Rolf et al. (2024) Esther Rolf, Konstantin Klemmer, Caleb Robinson, and Hannah Kerner. 2024. Mission Critical – Satellite Data is a Distinct Modality in Machine Learning. https://doi.org/10.48550/arXiv.2402.01444 arXiv:2402.01444 [cs].
@@
- Rolf et al. (2021) Esther Rolf, Jonathan Proctor, Tamma Carleton, Ian Bolliger, Vaishaal Shankar, Miyabi Ishihara, Benjamin Recht, and Solomon Hsiang. 2021. A generalizable and accessible approach to machine learning with global satellite imagery. Nature Communications 12, 1 (July 2021), 4392. https://doi.org/10.1038/s41467-021-24638-z
@@
- Rosenqvist et al. (2007) Ake Rosenqvist, Masanobu Shimada, Norimasa Ito, and Manabu Watanabe. 2007. ALOS PALSAR: A Pathfinder Mission for Global-Scale Monitoring of the Environment. IEEE Transactions on Geoscience and Remote Sensing 45, 11 (Nov. 2007), 3307–3316. https://doi.org/10.1109/TGRS.2007.901027 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Russakovsky et al. (2015) Olga Russakovsky, Jia Deng, Hao Su, Jonathan Krause, Sanjeev Satheesh, Sean Ma, Zhiheng Huang, Andrej Karpathy, Aditya Khosla, Michael Bernstein, Alexander C. Berg, and Li Fei-Fei. 2015. ImageNet Large Scale Visual Recognition Challenge. International Journal of Computer Vision 115, 3 (Dec. 2015), 211–252. https://doi.org/10.1007/s11263-015-0816-y
@@
- Rußwurm et al. (2024) Marc Rußwurm, Konstantin Klemmer, Esther Rolf, Robin Zbinden, and Devis Tuia. 2024. Geographic Location Encoding with Spherical Harmonics and Sinusoidal Representation Networks. http://arxiv.org/abs/2310.06743 arXiv:2310.06743 [cs].
@@
- Singh et al. (2022) Amanpreet Singh, Ronghang Hu, Vedanuj Goswami, Guillaume Couairon, Wojciech Galuba, Marcus Rohrbach, and Douwe Kiela. 2022. FLAVA: A Foundational Language And Vision Alignment Model. In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, New Orleans, LA, USA, 15617–15629. https://doi.org/10.1109/CVPR52688.2022.01519
@@
- Sitzmann et al. (2020) Vincent Sitzmann, Julien N. P. Martel, Alexander W. Bergman, David B. Lindell, and Gordon Wetzstein. 2020. Implicit Neural Representations with Periodic Activation Functions. https://doi.org/10.48550/arXiv.2006.09661 arXiv:2006.09661 [cs].
@@
- Solari et al. (2017) Lorenzo Solari, Andrea Ciampalini, Federico Raspini, Silvia Bianchini, Ivana Zinno, Manuela Bonano, Michele Manunta, Sandro Moretti, and Nicola Casagli. 2017. Combined Use of C- and X-Band SAR Data for Subsidence Monitoring in an Urban Area. Geosciences 7, 2 (June 2017), 21. https://doi.org/10.3390/geosciences7020021 Number: 2 Publisher: Multidisciplinary Digital Publishing Institute.
- Song et al. (2015) Shuran Song, Samuel P. Lichtenberg, and Jianxiong Xiao. 2015. SUN RGB-D: A RGB-D Scene Understanding Benchmark Suite. 567–576. https://openaccess.thecvf.com/content_cvpr_2015/html/Song_SUN_RGB-D_A_2015_CVPR_paper.html
@@
- Spoto et al. (2012) Francois Spoto, Omar Sy, Paolo Laberinti, Philippe Martimort, Valerie Fernandez, Olivier Colin, Bianca Hoersch, and Aime Meygret. 2012. Overview Of Sentinel-2. In 2012 IEEE International Geoscience and Remote Sensing Symposium. 1707–1710. https://doi.org/10.1109/IGARSS.2012.6351195 ISSN: 2153-7003.
@@
- Sumbul et al. (2019) Gencer Sumbul, Marcela Charfuelan, Begüm Demir, and Volker Markl. 2019. Bigearthnet: A Large-Scale Benchmark Archive for Remote Sensing Image Understanding. In IGARSS 2019 - 2019 IEEE International Geoscience and Remote Sensing Symposium. 5901–5904. https://doi.org/10.1109/IGARSS.2019.8900532 ISSN: 2153-7003.
@@
- Sun et al. (2023) Xian Sun, Peijin Wang, Wanxuan Lu, Zicong Zhu, Xiaonan Lu, Qibin He, Junxi Li, Xuee Rong, Zhujun Yang, Hao Chang, Qinglin He, Guang Yang, Ruiping Wang, Jiwen Lu, and Kun Fu. 2023. RingMo: A Remote Sensing Foundation Model With Masked Image Modeling. IEEE Transactions on Geoscience and Remote Sensing 61 (2023), 1–22. https://doi.org/10.1109/TGRS.2022.3194732 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Tachi et al. (1989) K. Tachi, K. Arai, and Y. Sato. 1989. Advanced microwave scanning radiometer (AMSR): requirements and preliminary design study. IEEE Transactions on Geoscience and Remote Sensing 27, 2 (March 1989), 177–183. https://doi.org/10.1109/36.20296 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Tang et al. (2024) Maofeng Tang, Andrei Cozma, Konstantinos Georgiou, and Hairong Qi. 2024. Cross-Scale MAE: A Tale of Multi-Scale Exploitation in Remote Sensing. https://doi.org/10.48550/arXiv.2401.15855 arXiv:2401.15855 [cs].
@@
- Thompson* (2015) Alan A. Thompson*. 2015. Overview of the RADARSAT Constellation Mission. Canadian Journal of Remote Sensing 41, 5 (Sept. 2015), 401–407. https://doi.org/10.1080/07038992.2015.1104633 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.1080/07038992.2015.1104633.
@@
- Torres et al. (2012) Ramon Torres, Paul Snoeij, Dirk Geudtner, David Bibby, Malcolm Davidson, Evert Attema, Pierre Potin, BjÖrn Rommen, Nicolas Floury, Mike Brown, Ignacio Navas Traver, Patrick Deghaye, Berthyl Duesmann, Betlem Rosich, Nuno Miranda, Claudio Bruno, Michelangelo L’Abbate, Renato Croci, Andrea Pietropaolo, Markus Huchler, and Friedhelm Rostan. 2012. GMES Sentinel-1 mission. Remote Sensing of Environment 120 (May 2012), 9–24. https://doi.org/10.1016/j.rse.2011.05.028
@@
- Tseng et al. (2024) Gabriel Tseng, Ruben Cartuyvels, Ivan Zvonkov, Mirali Purohit, David Rolnick, and Hannah Kerner. 2024. Lightweight, Pre-trained Transformers for Remote Sensing Timeseries. https://doi.org/10.48550/arXiv.2304.14065 arXiv:2304.14065 [cs].
@@
- Tsokas et al. (2022) Arsenios Tsokas, Maciej Rysz, Panos M. Pardalos, and Kathleen Dipple. 2022. SAR data applications in earth observation: An overview. Expert Systems with Applications 205 (Nov. 2022), 117342. https://doi.org/10.1016/j.eswa.2022.117342
- Van Horn et al. (2018) Grant Van Horn, Oisin Mac Aodha, Yang Song, Yin Cui, Chen Sun, Alex Shepard, Hartwig Adam, Pietro Perona, and Serge Belongie. 2018. The INaturalist Species Classification and Detection Dataset. 8769–8778. https://openaccess.thecvf.com/content_cvpr_2018/html/Van_Horn_The_INaturalist_Species_CVPR_2018_paper.html
@@
- Wang et al. (2023c) Di Wang, Qiming Zhang, Yufei Xu, Jing Zhang, Bo Du, Dacheng Tao, and Liangpei Zhang. 2023c. Advancing Plain Vision Transformer Toward Remote Sensing Foundation Model. IEEE Transactions on Geoscience and Remote Sensing 61 (2023), 1–15. https://doi.org/10.1109/TGRS.2022.3222818 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Wang et al. (2016) Qunming Wang, Wenzhong Shi, Zhongbin Li, and Peter M. Atkinson. 2016. Fusion of Sentinel-2 images. Remote Sensing of Environment 187 (Dec. 2016), 241–252. https://doi.org/10.1016/j.rse.2016.10.030
- Wang et al. (2024) Yi Wang, Conrad M. Albrecht, Nassim Ait Ali Braham, Chenying Liu, Zhitong Xiong, and Xiao Xiang Zhu. 2024. Decoupling Common and Unique Representations for Multimodal Self-supervised Learning. https://doi.org/10.48550/arXiv.2309.05300 arXiv:2309.05300 [cs].